Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by BTGmoderatorRO » Sun Dec 17, 2017 5:29 am
A certain team has 12 members, including Joey. A three-member relay team will be selected as follows: one of the 12 members is to be chosen at random to run fi…rst, one of the remaining 11 members is to be chosen at random to run second, and one of the remaining 10 members is to be chosen at random to run third. What is the probability that Joey will be chosen to run second or third?

(A) 1/1320
(B) 1/132
(C) 1/110
(D) 1/12
(E) 1/6

OA is E

I need an elaborate breakdown on why E is the correct answer. Can any expert assist here pls
Join the discussion
Source: — Problem Solving |

by [email protected] » Sun Dec 17, 2017 1:28 pm
Hi Roland2rule,

We're told that a certain team has 12 members, including Joey - and a three-member relay team will be selected as follows: one of the 12 members is to be chosen at random to run fi…rst, one of the remaining 11 members is to be chosen at random to run second, and one of the remaining 10 members is to be chosen at random to run third. We're asked for the probability that Joey will be chosen to run second OR third. Since there are two possible "successful" outcomes, we have to complete two calculations:

Probability of Joey running 2nd: = (Not Joey)(Joey)(Not Joey) = (11/12)(1/11)(10/10) = 1/12
Probability of Joey running 3rd: = (Not Joey)(Not Joey)(Joey) = (11/12)(10/11)(1/10) = 1/12

Probability of either of those options occurring = (1/12) + (1/12) = 2/12 = 1/6

Final Answer: E

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Scott@TargetTestPrep » Mon Sep 16, 2019 10:21 am
BTGmoderatorRO wrote:A certain team has 12 members, including Joey. A three-member relay team will be selected as follows: one of the 12 members is to be chosen at random to run fi…rst, one of the remaining 11 members is to be chosen at random to run second, and one of the remaining 10 members is to be chosen at random to run third. What is the probability that Joey will be chosen to run second or third?

(A) 1/1320
(B) 1/132
(C) 1/110
(D) 1/12
(E) 1/6

OA is E

I need an elaborate breakdown on why E is the correct answer. Can any expert assist here pls

The probability Joey will be chosen to run second is:

11/12 x 1/11 x 10/10 = 1/12

The probability Joey will be chosen to run third is:

11/12 x 10/11 x 1/10 = 1/12

Thus, the probability that he will be chosen to run second or third is:

1/12 + 1/12 = 2/12 = 1/6

Answer: E

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion